sadghostheroin:

Choose a system of equations with the same solution as the following system: 6x + 2y = −6 3x − 4y = −18

J4ke:

sorry but the equation editor is not really good any way do you know how to solve this two equations?

J4ke:

hint: multiply the first equation by 2 to get rid of y

sadghostheroin:

oh ok

J4ke:

....(1)..... 6x + 2y = −6 ....(2).....3x − 4y = −18 2*6x+2*2y=2*-6 =12x+4y=-12 the sum of (1)+(2) 12x+4y=-12 3x-4y=-18 15x=-30 x =-30/15

sadghostheroin:

oooooooo ok

sadghostheroin:

so that is the equation for both right?

J4ke:

after that you replace x in the equation with the real value to get y x = -30/15= -2

J4ke:

replace x here ....(2).....3x − 4y = −18

J4ke:

with -2

J4ke:

3(-2)-4y=-18 -6-4y=-18 -6+18=4y 12=4y y=12/4 y=3

sadghostheroin:

thank u so much tbh

J4ke:

do you really understand what i did or you want me to explain further more? :)

sadghostheroin:

i do i understand for like the first time :3 thank you